# Not All Sub-Neptune Exoplanets Have Magma Oceans

### v1

@@ Abstract
The evolution and structure of sub-Neptunes may be strongly influenced by interactions between the outer gaseous envelope of the planet and a surface magma ocean. However, given the wide variety of permissible interior structures of these planets, it is unclear whether conditions at the envelope-mantle boundary will always permit a molten silicate layer, or whether some sub-Neptunes might instead host a solid silicate surface. In this work, we use internal structure modeling to perform an extensive exploration of surface conditions within the sub-Neptune population across a range of bulk and atmospheric parameters. We find that a significant portion of the population may lack present-day magma oceans. In particular, planets with a high atmospheric mean molecular weight and large envelope mass fraction are likely to instead have a solid silicate surface, since the pressure at the envelope-mantle boundary is high enough that the silicates will be in solid post-perovskite phase. This result is particularly relevant given recent inferences of high-mean molecular weight atmospheres from JWST observations of several sub-Neptunes. We apply this approach to a number of sub-Neptunes with existing or upcoming JWST observations, and find that in almost all cases, a range of solutions exist which do not possess a present-day magma ocean. Our analysis provides critical context for interpreting sub-Neptunes and their atmospheres.
@@
software: SMILE (M. C. Nixon & N. Madhusudhan, 2021), NumPy (C. R. Harris et al., 2020), SciPy (P. Virtanen et al., 2020), Matplotlib (J. D. Hunter, 2007), Pandas (T. pandas development team, 2020)
@@ 1 Introduction
Modern exoplanet surveys have revealed an abundance of close-in planets with sizes ranging from 1-4 $R_{\oplus}$ (N. M. Batalha et al., 2013). This population exhibits a bimodal radius distribution, with a first peak at $\sim$1.3 $R_{\oplus}$, the second at $\sim$2.4 $R_{\oplus}$, and a minimum at $\sim$1.75 $R_{\oplus}$ (B. J. Fulton et al., 2017). Planets with radii << 1.75 $R_{\oplus}$ are often referred to as super-Earths, while those with radii >> 1.75 $R_{\oplus}$ are called sub-Neptunes. The bulk densities of sub-Neptunes indicate that they must possess some kind of volatile component, which could take the form of, e.g., a H/He envelope making up a few percent of the planet’s mass (E. D. Lopez & J. J. Fortney, 2014; L. A. Rogers, 2015; H. Chen & L. A. Rogers, 2016), or a water-rich envelope of up to $\sim 50\%$ of the planet’s mass (A. Léger et al., 2004; L. Zeng et al., 2019; R. Luque & E. Pallé, 2022). However, mass and radius alone are not enough to disentangle these possibilities (e.g., L. A. Rogers et al., 2011; J. G. Rogers et al., 2023). In order to understand the composition of these objects, we require atmospheric characterization, both through observational constraints on the upper atmosphere via spectroscopy (e.g., L. Kreidberg et al., 2014; E. M.-R. Kempton et al., 2023; B. Benneke et al., 2024; B. Davenport et al., 2025), and through detailed modeling of the key formation and evolutionary processes that shape these atmospheres throughout their lifetimes (e.g., J. E. Owen & Y. Wu, 2013; E. S. Kite et al., 2019; H. E. Schlichting & E. D. Young, 2022).
@@
One of the key processes that shapes the evolution of sub-Neptunes arises from the fact that their gaseous envelopes are expected to be in contact with a long-lived magma ocean (E. S. Kite et al., 2019). Interactions between the magma and the envelope have been invoked to explain the steep decline in the frequency of sub-Neptunes with radii above 3 $R_{\oplus}$ (B. J. Fulton et al., 2017). This occurs because magma oceans will limit growth due to the solubility of hydrogen in magma at high pressures (E. S. Kite et al., 2019). Furthermore, the interaction between a magma ocean and a H2-rich envelope can affect atmospheric composition and structure (e.g., H. E. Schlichting & E. D. Young, 2022; W. Misener et al., 2023; C. Seo et al., 2024), potentially leading to observable atmospheric signatures (e.g., F. E. Rigby et al., 2024; O. Shorttle et al., 2024; A. Werlen et al., 2025).
@@
Although magma oceans may be crucial in affecting the atmospheric composition and structure of sub-Neptunes, it is currently unclear whether the silicate layer of a given sub-Neptune is expected to be molten at all times. The thermodynamic conditions at the envelope-mantle boundary will be governed by numerous parameters, including the planet mass, temperature profile and atmospheric composition. Although it has been noted in some cases that certain sub-Neptunes may not currently host magma oceans (F. E. Rigby et al., 2024), there is very little work exploring the parameter space for which magma oceans are expected to persist or solidify. Additionally, molecular dynamics simulations have shown that for sub-Neptunes, hydrogen and water may be distributed in the core and mantle as well as the gaseous envelope (L. Stixrude & B. Karki, 2005; H. Luo et al., 2024). However, in order to determine whether the thermodynamic conditions of these simulations are relevant to sub-Neptune interiors, we must first evaluate the range of possible temperatures and pressures that could occur at the envelope-mantle boundary for planets with different bulk and atmospheric properties.
@@
JWST has revealed a wide range of atmospheric compositions within the sub-Neptune population (e.g. N. Madhusudhan et al., 2023; B. Benneke et al., 2024; C. Piaulet-Ghorayeb et al., 2024; B. Davenport et al., 2025). JWST/MIRI observations of sub-Neptune GJ 1214 b (E. M.-R. Kempton et al., 2023) found a metal rich upper atmosphere, with P. Gao et al. (2023) finding a $\gtrsim$ 300x solar metallicity, implying a mean molecular weight (MMW) of $\gtrsim$ 9.8 g/mol. Additional observations using JWST/NIRSpec also point to a high-MMW atmosphere (E. Schlawin et al., 2024; K. Ohno et al., 2025). Internal structure modeling informed by the atmospheric constraints from E. M.-R. Kempton et al. (2023) and P. Gao et al. (2023) suggests a high envelope mass fraction ($x_{\rm env}$), resulting in high pressures at the envelope-mantle boundary that could push the rocky surface into a solid phase (M. C. Nixon et al., 2024). GJ 1214 b is not the only sub-Neptune observed to have a high-MMW atmosphere, with JWST observations suggesting a MMW >5>5\,g/mol for TOI-270 d (B. Benneke et al., 2024), GJ 9827 d (C. Piaulet-Ghorayeb et al., 2024) and TOI-836 c (N. L. Wallack et al., 2024). These findings motivate a wider parameter exploration to determine the prevalence of magma oceans on sub-Neptunes with a range of bulk and atmospheric properties. Due to the important role which magma oceans play in both atmospheric composition and planetary structure, it is vital that we explore the possibility of their absence in a subset of sub-Neptunes.
@@ 2.1 Internal Structure Model
We use the same internal structure model as in M. C. Nixon et al. (2024), which is an updated version of the model initially presented in M. C. Nixon & N. Madhusudhan (2021), titled Structural Model of Internal Layers of Exoplanets (SMILE111github.com/mcnixon/smile). This model solves the mass continuity and hydrostatic equilibrium equations using numerical integration and assuming spherical symmetry. Planets consist of an iron core, silicate mantle, and a mixed H/He/H2O envelope. We solve for the planetary radius $R_{p}$ for an input mass $M_{p}$, mass fractions $x_{i}$ = $M_{i}$$M_{p}$ of each component, and envelope MMW. The mass fractions of H2O and H/He are calculated from the MMW (see equation 1, M. C. Nixon et al., 2024). Throughout this work, we assume an Earth-like iron-to-silicate ratio (1/3 iron, 2/3 silicates by mass).
@@
SMILE uses an isothermal equation of state (EOS) for iron, adopted from S. Seager et al. (2007). Thermal expansion is negligible for iron at high pressures (A. R. Howe et al., 2014). The silicate EOS prescription is temperature-dependent, following the phase diagram shown in C. Huang et al. (2022), and consists of three phases: bridgmanite (A. R. Oganov & S. Ono, 2004), post-perovskite (T. Sakai et al., 2016), and liquid MgSiO3 (A. S. Wolf & D. J. Bower, 2018). The silicate phase diagram can be seen in Figure 1.
@@
A temperature-dependent EOS is also used for the outer gaseous envelope, as the temperature profile of the envelope can significantly affect the radii of sub-Neptunes (M. C. Nixon & N. Madhusudhan, 2021). The EOS for H2O is adopted from M. C. Nixon & N. Madhusudhan (2021) and was constructed from a range of sources, covering the full pressure-temperature (PPTT) space which is relevant to sub-Neptunes (E. E. Salpeter & H. S. Zapolsky, 1967; Y. Fei et al., 1993; W. Wagner & A. Pruß, 2002; R. Feistel & W. Wagner, 2006; S. Seager et al., 2007; M. French et al., 2009; S. Klotz et al., 2017; B. Journaux et al., 2020). The H/He EOS is adopted from G. Chabrier et al. (2019) and assumes a helium mass fraction Y = 0.275. Our implementation of the SMILE model assumes a fully mixed H/He/H2O envelope, as predicted by simulations of sub-Neptunes with equilibrium temperatures down to $\sim$300 K (A. Gupta et al., 2025). An ideal mixing approximation is used to construct the mixed EOS, which has been shown to be appropriate for these components (F. Soubiran & B. Militzer, 2015).
@@
The photospheric temperature ($T_{0}$) and pressure ($P_{0}$) are also required input parameters. While we allow $T_{0}$ to vary as part of our grid, we fix $P_{0}$ to a value of 100 Pa, or $10^{-3}$ bar. This is a typical photospheric pressure for an sub-Neptune atmosphere in transit geometry (e.g., A. A. A. Piette & N. Madhusudhan, 2020). We adopt PPTT profiles consisting of an isotherm down to some radiative-convective boundary (determined by a pressure $P_{\rm rc}$), at which point the model switches to follow an adiabatic profile. $P_{\rm rc}$ serves as an additional input to the model. Higher $P_{\rm rc}$ values extend the isothermal region deeper into the envelope, leading to a cooler temperature profile overall. This mechanism serves as a proxy for planetary age, where lower $P_{\rm rc}$ values correspond to younger and hotter planets, a relationship which has been demonstrated for evolutionary models (J. G. Rogers, 2025). By varying $P_{\rm rc}$, we can draw qualitative conclusions regarding the time evolution of these planets. Although this is a relatively simple prescription for the PPTT profile, it has been shown to be adequate for interior structure modeling (M. C. Nixon & N. Madhusudhan, 2021; J. G. Rogers, 2025).
@@ 2.2 Model Grid
The minimum $M_{p}$ of 4 $M_{\oplus}$ is an approximate lower bound for sub-Neptunes with substantial gaseous envelopes (L. Parc et al., 2024) while the maximum planetary mass of 20 $M_{\oplus}$ is informed by the fact that Neptune itself is only 17 $M_{\oplus}$, alongside constraints from Kepler on the mass of planetary cores (J. E. Owen & Y. Wu, 2013). The lower limit for $T_{0}$ of 300 K is informed by the fact that lower temperatures than this can lead to condensation of H2O and possible stratification of the envelope (B. Benneke et al., 2024), which we do not consider in this study. The upper limit is comparable to the hottest sub-Neptune whose upper atmosphere has been constrained with JWST spectroscopy, TOI-421 b (B. Davenport et al., 2025). We note that very hot sub-Neptunes ($T_{\rm eq}\gg 1000\,$K) are likely to lose their envelopes due to extreme stellar irradiation (J. E. Owen & Y. Wu, 2017). $P_{\rm rc}$ ranges from 1 to 100 bar on a logarithmic scale, informed by self-consistent atmospheric models of sub-Neptunes (e.g., A. A. A. Piette & N. Madhusudhan, 2020). We vary $x_{\rm env}$ from 0.1%–50%, with 0.1% serving as a minimum for planets likely to retain gaseous envelopes (J. E. Owen & Y. Wu, 2017) and 50% as a likely maximum limit for ice accretion highlighted in K. Lodders (2003) and L. Zeng et al. (2019). The MMW ranges from 2.35 g/mol (pure H/He) to 18.02 g/mol (pure H2O).
@@ 3.1 Initial Case Study: GJ 1214 b
Figure 1: Pressure-temperature profiles of GJ 1214 b with different compositions ($300\times$ and $500\times$ solar metallicity) and haze properties (maximally reflective hazes and tholin hazes), from M. C. Nixon et al. (2024). The profiles are overlaid on the phase diagram of MgSiO3, showing the solid-liquid and bridgmanite-post-perovskite (Brg-PPv) boundaries. Circles indicate the pressure and temperature at the envelope-mantle boundary for each model. In all cases, the boundary occurs in the solid phase. The dotted orange line shows the best-fit isothermal-adiabatic profile (i.e., the parameterization used in this work) to the $300\times$ solar, maximally reflective haze model.
As an initial study, we consider the planet GJ 1214 b (D. Charbonneau et al., 2009), one of the most extensively studied sub-Neptunes. Observing campaigns from the ground as well as using HST and JWST have determined that this planet hosts a hazy, high-MMW atmosphere (J. L. Bean et al., 2010, 2011; J.-M. Désert et al., 2011; Z. K. Berta et al., 2012; E. J. W. de Mooij et al., 2012; J. K. Teske et al., 2013; L. Kreidberg et al., 2014; E. M.-R. Kempton et al., 2023; P. Gao et al., 2023; E. Schlawin et al., 2024; K. Ohno et al., 2025). These results, alongside self-consistent radiative transfer calculations to predict the atmospheric structure, were used to inform internal structure models to infer the possible bulk composition of the planet (M. C. Nixon et al., 2024), revealing a substantial gaseous envelope with pressures and temperatures at the envelope-mantle boundary that correspond to a solid (i.e., non-magma) mantle surface. Our investigation of GJ 1214 b serves two main functions: (1) to further explore the extent to which the set of possible internal structures of GJ 1214 b permit a solid envelope-mantle boundary, and (2) to test the validity of the simpler isothermal-adiabatic PPTT profile assumption that we use in the remainder of this study.
@@
We use this example case to validate our simplified PPTT profile prescription used in the following section. For each of the four self-consistent PPTT profiles considered, we find a corresponding isothermal-adiabatic profile that closely matches the deep atmospheric temperature. In order to do this, we fix $T_{0}$ at the planet’s equilibrium temperature (596 K, R. Cloutier et al., 2021) and optimize $P_{\rm rc}$ by conducting a bisection search in log-space until $\Delta T/T<1\%$ for all pressures in the deep atmosphere (P>100P>100\,bar). We verify that the planet radius as well as the temperature and pressure at the envelope-mantle boundary are similar to the values found when using the self-consistent PPTT profile. Having confirmed this, we justify using the simplified isothermal-adiabatic PPTT profiles for our large model grid.

### v2

@@ Abstract
The evolution and structure of sub-Neptunes may be strongly influenced by interactions between the outer gaseous envelope of the planet and a surface magma ocean. However, given the wide variety of permissible interior structures of these planets, it is unclear whether conditions at the envelope-mantle boundary will always permit a molten silicate layer, or whether some sub-Neptunes might instead host a solid silicate surface. In this work, we use internal structure modeling to perform an extensive exploration of surface conditions within the sub-Neptune population across a range of bulk and atmospheric parameters. We find that a significant portion of the population may lack present-day magma oceans. In particular, planets with a high atmospheric mean molecular weight and large envelope mass fraction are likely to instead have a solid silicate surface, since the pressure at the envelope-mantle boundary is high enough that the silicates will be in solid postperovskite phase. This result is particularly relevant given recent inferences of high-mean molecular weight atmospheres from JWST observations of several sub-Neptunes. We apply this approach to a number of sub-Neptunes with existing or upcoming JWST observations, and find that in almost all cases, a range of solutions exist which do not possess a present-day magma ocean. Our analysis provides critical context for interpreting sub-Neptunes and their atmospheres.
@@
software: SMILE (Nixon_Madhusudhan_2021), NumPy (Harris2020), SciPy (Virtanen2020), Matplotlib (Hunter2007), Pandas (pandas)
@@ 1 Introduction
Modern exoplanet surveys have revealed an abundance of close-in planets with sizes ranging from 1-4 $R_{\oplus}$ (Batalha_2013). This population exhibits a bimodal radius distribution, with a first peak at $\sim$1.3 $R_{\oplus}$, the second at $\sim$2.4 $R_{\oplus}$, and a minimum at $\sim$1.75 $R_{\oplus}$ (Fulton_2017). Planets with radii << 1.75 $R_{\oplus}$ are often referred to as super-Earths, while those with radii >> 1.75 $R_{\oplus}$ are called sub-Neptunes. The bulk densities of sub-Neptunes indicate that they must possess some kind of volatile component, which could take the form of, e.g., a H/He envelope making up a few percent of the planet’s mass (Lopez_2014; Rogers_2015; Chen_2016), or a water-rich envelope of up to $\sim 50\%$ of the planet’s mass (Leger_2004; Zeng_2019; Luque_2022). However, mass and radius alone are not enough to disentangle these possibilities (e.g., Rogers_2011; Rogers_2023). In order to understand the composition of these objects, we require atmospheric characterization, both through observational constraints on the upper atmosphere via spectroscopy (e.g., Kreidberg_2014; Kempton_2023; Benneke_2024; Davenport_2025), and through detailed modeling of the key formation and evolutionary processes that shape these atmospheres throughout their lifetimes (e.g., Owen_2013; Kite_2019; Schlichting_2022).
@@
One of the key processes that shapes the evolution of sub-Neptunes arises from the fact that their gaseous envelopes are expected to be in contact with a long-lived magma ocean (Kite_2019). Interactions between the magma and the envelope have been invoked to explain the steep decline in the frequency of sub-Neptunes with radii above 3 $R_{\oplus}$ (Fulton_2017). This occurs because magma oceans will limit growth due to the solubility of hydrogen in magma at high pressures (Kite_2019). Furthermore, the interaction between a magma ocean and a H2-rich envelope can affect atmospheric composition and structure (e.g., Schlichting_2022; Misener_2023; Seo_2024), potentially leading to observable atmospheric signatures (e.g., Rigby_2024; Shorttle_2024; Werlen_2025).
@@
Although magma oceans may be crucial in affecting the atmospheric composition and structure of sub-Neptunes, it is currently unclear whether the silicate layer of a given sub-Neptune is expected to be molten at all times. The thermodynamic conditions at the envelope-mantle boundary will be governed by numerous parameters, including the planet mass, temperature profile and atmospheric composition. Although it has been noted in some cases that certain sub-Neptunes may not currently host magma oceans (Rigby_2024), there is very little work exploring the parameter space for which magma oceans are expected to persist or solidify. Additionally, molecular dynamics simulations have shown that for sub-Neptunes, hydrogen and water may be distributed in the core and mantle as well as the gaseous envelope (Stixrude2005; Luo2024). However, in order to determine whether the thermodynamic conditions of these simulations are relevant to sub-Neptune interiors, we must first evaluate the range of possible temperatures and pressures that could occur at the envelope-mantle boundary for planets with different bulk and atmospheric properties.
@@
JWST has revealed a wide range of atmospheric compositions within the sub-Neptune population (e.g. Madhusudhan_2023; Benneke_2024; Piaulet-Ghorayeb_2024; Davenport_2025). JWST/MIRI observations of sub-Neptune GJ 1214 b (Kempton_2023) found a metal rich upper atmosphere, with Gao_2023 finding a $\gtrsim$ 300x solar metallicity, implying a mean molecular weight (MMW) of $\gtrsim$ 9.8 g/mol. Additional observations using JWST/NIRSpec also point to a high-MMW atmosphere (Schlawin2024; Ohno2025). Internal structure modeling informed by the atmospheric constraints from Kempton_2023 and Gao_2023 suggests a high envelope mass fraction ($x_{\rm env}$), resulting in high pressures at the envelope-mantle boundary that could push the rocky surface into a solid phase (Nixon_2024). GJ 1214 b is not the only sub-Neptune observed to have a high-MMW atmosphere, with JWST observations suggesting a MMW >5>5\,g/mol for TOI-270 d (Benneke_2024), GJ 9827 d (Piaulet-Ghorayeb_2024) and TOI-836 c (Wallack2024). These findings motivate a wider parameter exploration to determine the prevalence of magma oceans on sub-Neptunes with a range of bulk and atmospheric properties. Due to the important role which magma oceans play in both atmospheric composition and planetary structure, it is vital that we explore the possibility of their absence in a subset of sub-Neptunes.
@@ 2.1 Internal Structure Model
We use the same internal structure model as in Nixon_2024, which is an updated version of the model initially presented in Nixon_Madhusudhan_2021, titled Structural Model of Internal Layers of Exoplanets (SMILE111github.com/mcnixon/smile). This model solves the mass continuity and hydrostatic equilibrium equations using numerical integration and assuming spherical symmetry. Planets consist of an iron core, silicate mantle, and a mixed H/He/H2O envelope. We solve for the planetary radius $R_{p}$ for an input mass $M_{p}$, mass fractions $x_{i}$ = $M_{i}$$M_{p}$ of each component, and envelope MMW. The mass fractions of H2O and H/He are calculated from the MMW (see equation 1, Nixon_2024). Throughout this work, we assume an Earth-like iron-to-silicate ratio (1/3 iron, 2/3 silicates by mass).
@@
SMILE uses an isothermal equation of state (EOS) for iron, adopted from Seager_2007. Thermal expansion is negligible for iron at high pressures (Howe_2014). The silicate EOS prescription is temperature-dependent, following the phase diagram shown in Huang_2022, and consists of three phases: bridgmanite (Oganov_2004), postperovskite (Sakai_2016), and liquid MgSiO3 (Wolf_2018). The silicate phase diagram can be seen in Figure 1.
@@
A temperature-dependent EOS is also used for the outer gaseous envelope, as the temperature profile of the envelope can significantly affect the radii of sub-Neptunes (Nixon_Madhusudhan_2021). The EOS for H2O is adopted from Nixon_Madhusudhan_2021 and was constructed from a range of sources, covering the full pressure-temperature (PPTT) space which is relevant to sub-Neptunes (Salpeter_1967; Fei_1993; Wagner_2002; Feistel_2006; Seager_2007; French_2009; Klotz_2017; Journaux_2020). The H/He EOS is adopted from Chabrier_2019 and assumes a helium mass fraction Y = 0.275. Our implementation of the SMILE model assumes a fully mixed H/He/H2O envelope, as predicted by simulations of sub-Neptunes with equilibrium temperatures down to $\sim$300 K (Gupta2025). An ideal mixing approximation is used to construct the mixed EOS, which has been shown to be appropriate for these components (Soubiran_2015).
@@
The photospheric temperature ($T_{0}$) and pressure ($P_{0}$) are also required input parameters. While we allow $T_{0}$ to vary as part of our grid, we fix $P_{0}$ to a value of 100 Pa, or $10^{-3}$ bar. This is a typical photospheric pressure for an sub-Neptune atmosphere in transit geometry (e.g., Piette_2020). We adopt PPTT profiles consisting of an isotherm down to some radiative-convective boundary (determined by a pressure $P_{\rm rc}$), at which point the model switches to follow an adiabatic profile. $P_{\rm rc}$ serves as an additional input to the model. Higher $P_{\rm rc}$ values extend the isothermal region deeper into the envelope, leading to a cooler temperature profile overall. This mechanism serves as a proxy for planetary age, where lower $P_{\rm rc}$ values correspond to younger and hotter planets, a relationship which has been demonstrated for evolutionary models (Rogers2025). By varying $P_{\rm rc}$, we can draw qualitative conclusions regarding the time evolution of these planets. Although this is a relatively simple prescription for the PPTT profile, it has been shown to be adequate for interior structure modeling (Nixon_Madhusudhan_2021; Rogers2025).
@@ 2.2 Model Grid
The minimum $M_{p}$ of 4 $M_{\oplus}$ is an approximate lower bound for sub-Neptunes with substantial gaseous envelopes (Parc_2024) while the maximum planetary mass of 20 $M_{\oplus}$ is informed by the fact that Neptune itself is only 17 $M_{\oplus}$, alongside constraints from Kepler on the mass of planetary cores (Owen_2013). The lower limit for $T_{0}$ of 300 K is informed by the fact that lower temperatures than this can lead to condensation of H2O and possible stratification of the envelope (Benneke_2024), which we do not consider in this study. The upper limit is comparable to the hottest sub-Neptune whose upper atmosphere has been constrained with JWST spectroscopy, TOI-421 b (Davenport_2025). We note that very hot sub-Neptunes ($T_{\rm eq}\gg 1000\,$K) are likely to lose their envelopes due to extreme stellar irradiation (Owen_2017). $P_{\rm rc}$ ranges from 1 to 100 bar on a logarithmic scale, informed by self-consistent atmospheric models of sub-Neptunes (e.g., Piette_2020). We vary $x_{\rm env}$ from 0.1%–50%, with 0.1% serving as a minimum for planets likely to retain gaseous envelopes (Owen_2017) and 50% as a likely maximum limit for ice accretion highlighted in Lodders_2003 and Zeng_2019. The MMW ranges from 2.35 g/mol (pure H/He) to 18.02 g/mol (pure H2O).
@@ 3.1 Initial Case Study: GJ 1214 b
Figure 1: Pressure-temperature profiles of GJ 1214 b with different compositions ($300\times$ and $500\times$ solar metallicity) and haze properties (maximally reflective hazes and tholin hazes), from Nixon_2024. The profiles are overlaid on the phase diagram of MgSiO3, showing the solid-liquid and bridgmanite-postperovskite (Brg-PPv) boundaries. Circles indicate the pressure and temperature at the envelope-mantle boundary for each model. In all cases, the boundary occurs in the solid phase. The dotted orange line shows the best-fit isothermal-adiabatic profile (i.e., the parameterization used in this work) to the $300\times$ solar, maximally reflective haze model.
As an initial study, we consider the planet GJ 1214 b (Charbonneau2009), one of the most extensively studied sub-Neptunes. Observing campaigns from the ground as well as using HST and JWST have determined that this planet hosts a hazy, high-MMW atmosphere (Bean2010; Bean2011; Desert2011; Berta2012; deMooij2012; Teske2013; Kreidberg_2014; Kempton_2023; Gao_2023; Schlawin2024; Ohno2025). These results, alongside self-consistent radiative transfer calculations to predict the atmospheric structure, were used to inform internal structure models to infer the possible bulk composition of the planet (Nixon_2024), revealing a substantial gaseous envelope with pressures and temperatures at the envelope-mantle boundary that correspond to a solid (i.e., non-magma) mantle surface. Our investigation of GJ 1214 b serves two main functions: (1) to further explore the extent to which the set of possible internal structures of GJ 1214 b permit a solid envelope-mantle boundary, and (2) to test the validity of the simpler isothermal-adiabatic PPTT profile assumption that we use in the remainder of this study.
@@
We use this example case to validate our simplified PPTT profile prescription used in the following section. For each of the four self-consistent PPTT profiles considered, we find a corresponding isothermal-adiabatic profile that closely matches the deep atmospheric temperature. In order to do this, we fix $T_{0}$ at the planet’s equilibrium temperature (596 K, Cloutier2021) and optimize $P_{\rm rc}$ by conducting a bisection search in log-space until $\Delta T/T<1\%$ for all pressures in the deep atmosphere (P>100P>100\,bar). We verify that the planet radius as well as the temperature and pressure at the envelope-mantle boundary are similar to the values found when using the self-consistent PPTT profile. Having confirmed this, we justify using the simplified isothermal-adiabatic PPTT profiles for our large model grid.